# Diffusion Cocktail: Fused Generation from Diffusion Models

### v1

@@ 3.3 Ditail for Model Style Transfer
6:${\boldsymbol{c}}\leftarrow\alpha\,{\boldsymbol{e}}_{pos}-\beta\,{\boldsymbol{e}}_{neg}$
@@
In Line 2-3 of Alg. 3, we first generate a guidance condition vector ${\boldsymbol{c}}$, an interpolation between the positive and negative prompt embedding using two hyper-parameters $\alpha,\beta$. $\alpha$ and $\beta$ are extra knobs that allow the user to control the strength of the guidance for content injection. As shown in Fig 6, a larger $\alpha$ value increases the impact of the source image content, mainly generated by interpreting the positive prompt. Conversely, a larger $\beta$ value does not always yield a significant outcome, as it depends on the semantics of the negative prompt. If the negative prompt is relevant to the content/structure, then $\beta$ will change the resulting content guidance accordingly (Sec. 4 Fig 10). We note that the guidance of $\alpha,\beta$ solely could control the content information: i.e., we use the style DM to generate the output using the guided input latent without injection. But only when combined with the Ditail content injection can we get the best results (more details in Sec. 4).
@@
Figure 6: Visualization of noisy latent after DM inversion with different scaling factor $\alpha$. A larger $\alpha$ leads to a stronger structure prior embedded in the noisy latent, resulting in a more structure-preserved starting point in the reverse sampling process.
@@
We have two extra inputs for applying Ditail on the image style transfer task. Firstly, the user may provide positive and negative prompts to describe the image to help the inversion process. We have also tested empty prompts, prompts generated from image captioning, and irrelevant prompts in Sec. 4 (Fig. 9), which verify that both empty and captioning prompts could generate promising results. Secondly, the user needs to specify a DM for diffusion inversion. Ideally, a DM matching the content image style is desirable. Empirically, the dependence on the inversion model is not strict. Results in Fig. 7 show that the generation results using different inversion DMs are similar for real content images. Compared to the novel generation method as described previously, we may want to apply the style transfer algorithm instead for the following two reasons: 1) the generating DM or the latent representations are not available, and 2) we want to adjust the guidance parameters $\alpha,\beta$ to strengthen or weaken the content transfer.
@@
We note this is a novel style transfer setting as we use DMs to set the styles of interest. This allows us to benefit from the extensive collection of fine-tuned DMs with drastically different styles. If a certain style not included in the available DMs is desirable, we can fine-tune a DM model on images of that style. Compared to the ordinary style transfer, this is equivalent to defining the target style using a set of images instead of only one image. On the one hand, this gives us more flexibility to define the desirable styles, and we may even mix images of different styles into a new style. On the other hand, the set of images tends to provide more stable and comprehensive style information compared to a single image. However, such advantages come at the cost of collecting similar-style images and fine-tuning the DM. But once the target style DM gets trained, we can apply it to transfer styles on numerous content images with Ditail, so the cost goes to zero asymptotically.
Figure 7: Comparison of different inversion models.
@@ 4 Experiments
We apply Ditail to multiple tasks and conduct qualitative and quantitative evaluations. Sec. 4.1 performs the proposed novel generation process on generated images and compares the style transfer results with various state-of-the-art methods. Sec. 4.2 ablates the effect of the proposed scaling factors and content injection in terms of fidelity, structure, and style. Sec. 4.3 dives into real image editing empowered by the Ditail framework. More qualitative and quantitative results can be found in the appendix.
Scaling Injection CLIP $\uparrow$ DINO $\downarrow$ $\text{FID}_{\text{tgt}}\downarrow$
$\alpha=1,\beta=0$ 0.1602 0.0849 173.98
$\alpha=2,\beta=0$ 0.1589 0.0760 173.80
$\alpha=4,\beta=0$ 0.1576 0.0508 188.61
$\alpha=8,\beta=0$ 0.1596 0.0352 214.32
$\alpha=1,\beta=1$ 0.1596 0.0571 186.46
$\alpha=2,\beta=1$ 0.1590 0.0647 179.71
$\alpha=4,\beta=1$ 0.1576 0.0562 182.86
$\alpha=8,\beta=1$ 0.1592 0.0367 212.20
$\alpha=1,\beta=0$checkmark 0.1646 0.0425 196.41
$\alpha=2,\beta=0$checkmark 0.1597 0.0354 194.02
$\alpha=4,\beta=0$checkmark 0.1563 0.0267 209.23
$\alpha=8,\beta=0$checkmark 0.1581 0.0245 228.05
$\alpha=1,\beta=1$checkmark 0.1580 0.0290 204.56
$\alpha=2,\beta=1$checkmark 0.1578 0.0320 197.71
$\alpha=4,\beta=1$checkmark 0.1563 0.0279 203.23
$\alpha=8,\beta=1$checkmark 0.1574 0.0249 226.94
Table 1: Quantitative results for model style transfer on generated images from the GEMRec-18K [7] dataset. We manually pick 5 models with distinctive styles and conduct style transfer on 90 textual prompts. The metrics are averaged over all 20 possible style transfer combinations. The mean CLIP cosine similarity  [9, 26] on source images is 0.1589, and the mean FID [10] distance between source and target images is 227.27. The best-performed metric for each group are shown in bold.
@@ 4.1 Results on Generated Images
Dataset and Model Source. Utilizing Ditial for style transfer requires DM checkpoints with various style specializations. The GEMRec-18K [7] dataset contains 18 thousand images generated by 200 text-to-image DMs fine-tuned on Stable Diffusion  [28], where the DM checkpoints were collected from Civitai. Each DM generates 90 images, and the prompts are divided into 12 categories.
Figure 8: Qualitative comparison of style transfer results on generated images drawn from GEMRec-18K [7]. The ControlNet [40] results are produced by its soft edge pre-processing and then fed into a generic img-to-img translation framework.
@@
We manually pick five DM checkpoints with distinctive style specialization and perform any-to-any style transfer using the Ditail framework. We use the images $\{\mathcal{I}^{(1)}_{i}\}_{i=1:90}$ generated by one source DM $\theta^{(1)}$ as the content image and stylize it with another target DM $\theta^{(2)}$, resulting in $\{\mathcal{I}^{(1)\rightarrow(2)}_{i}\}_{i=1:90}$. We compare the qualitative results with other style transfer baselines [21, 40, 15, 41] in Fig. 8. The content and style images are generated by $\theta^{(1)}$ and $\theta^{(2)}$ on the same prompt, respectively. Notably, performing style transfer with Ditail does not require a style image, whereas it is a necessity for conventional style transfer approaches. One key observation here is that the task setting for fine-tuning with one or a few style examples is limited, as it may not secure a stable style learning. Failing to capture certain styles will generally cause a near-unmodified output image, such as the car example for InST [41] and the boy example for the others. It is undeniable that the term “style” has been loosely defined, and most style transfer works tend to specialize in the styles elicited from artistic images. However, we want to expand our scope to a more generalizable setting, which involves practical transfer settings such as “realistic $\rightarrow$ anime” and “watercolor $\rightarrow$ illustration”, or even the nuance style that a single style image or plain text guidance can hardly capture. In such cases, numerous DM checkpoints with stable styles can serve as a plug-and-play choice and exhibit reasonable performance.
4.2 Ablation Study
@@
Quantitative Results. To quantify the change in image fidelity, structure, and style, we use three metrics, respectively. We use the cosine similarity in CLIP-Score [9, 26] to measure the compliance between prompt and image; we use the DINO self-similarity loss [1, 36] to measure the preservation of structure, which essentially compares the self-attention maps between source and target images; we use Fréchet Inception Distance (FID) [10] to the images generated by the target model to measure the closeness of style. We observe the following: First, all CLIP scores do not deviate too much from the source image baseline, and the transferred images from all groups move closer to the distribution of target images. Second, enlarging the scaling factor $\alpha$ tends to degrade the fidelity of generated images, preserve more structure from the source images, and push the transferred images further away from the target distribution. Third, introducing $\beta$ when $\alpha$ is relatively small helps preserve the structure. Forth, the best stylized results are obtained with $\alpha=2$. Lastly, similar to the conclusions from qualitative results, applying content injection alone already provides a strong baseline for structure preservation.
Figure 9: Comparisons of stylized real image editing with various prompts. We set $\omega=7.5$ (left) and $15$ (right) for each image pair.
Figure 10: Ablation on content injection and scaling $(\alpha,\beta)$.
Qualitative Results. As demonstrated in Fig. 10, we can see that solely using scaling factors or content injection can preserve the content structure to some extent. By taking a closer look, we draw the following conclusions: First, scaling $\alpha$ leads to better structure preservation, which complies with the intuition provided by Fig. 6. This plug-and-play technique does not introduce extra computational costs and might be applicable to other tasks that needs structure preservation. Second, introducing $\beta$ is beneficial in polishing the subtle details of the image (e.g., blurriness, lighting and shadow, coloring, etc.) and thereby enhances the overall quality. However, the effect of $\beta$ is quite case-sensitive, as it may not exhibit a quality boost and may potentially cause unnatural coloring of the images, such as the corgi example with $(\alpha,\beta)=(2,2)$. Lastly, applying content injection gives a remarkable structure-preserving performance. Applying both tricks together enables us to keep the content unchanged while shifting to the target style seamlessly.
@@ 4.3 Results on Real Images
Scaling Injection CLIP $\uparrow$ DINO $\downarrow$ $\text{FID}_{\text{tgt}}\downarrow$
$\alpha=1$ 0.0701 0.0920 109.33
$\alpha=2$ 0.0694 0.0852 109.53
$\alpha=4$ 0.0792 0.0609 118.37
$\alpha=8$ 0.0870 0.0478 135.16
$\alpha=1$checkmark 0.0761 0.0517 129.84
$\alpha=2$checkmark 0.0772 0.0449 125.32
$\alpha=4$checkmark 0.0882 0.0322 132.39
$\alpha=8$checkmark 0.0950 0.0297 147.24
Table 2: Quantitative results for model style transfer on 300 real images and captions sampled from the COCO Captions [2] dataset and transfer to 5 stylized DMs. We also use the stylized models to generate pseudo-target images with captions. The mean CLIP cosine similarity [9, 26] on source images is 0.0907, and the mean FID [10] distance between source and target images is 155.43. The best-performed metrics are shown in bold.
Data Source of Real Images. We randomly sample 300 images (with one caption per image) from the COCO Captions [2] dataset. Some demonstration images are drawn from the LAION-5B [32] dataset with textual prompts generated by the captioner model from BLIP [19].
Style Transfer. We perform style transfer on all five stylized models and compare the quantitative results using the aforementioned metrics. The results are presented in Tab. 2. One minor difference here is that we do not have negative prompts from the collected data, plus the benefit of introducing $\beta$ is not generalizable enough according to Tab. 1, so we skip the comparison on $\beta$ for this section. Overall, the quantitative patterns remain largely the same compared to the results on generated images, except that the best-performed CLIP cosine similarity is dominated by the groups with the largest $\alpha$. We hypothesize that this is because the caption of real images does not necessarily capture all the components of the images, whereas the generated images are produced from the given prompts. With a larger $\alpha$, the CLIP scores may still fluctuate for generated images but increase in the case of real images.
Stylized Image Editing. The Ditail framework enables stylized real-world image editing as well. Some qualitative results over various prompt types are shown in Fig. 9. The patterns we observed are as follows: First, null prompts hardly exhibit editing power and generally result in slightly stylized output. Second, a larger guidance scale $\omega$ [11] allows the target models to be more creative. Third, given that we enforce a strong structure emphasis through condition scaling and content injection, the editing prompts that change the number of objects usually give unsatisfactory results, whereas the semantic class of the objects might be modified effectively, even for completely unrelated prompts (e.g., car $\rightarrow$ boar, man $\rightarrow$ chicken, etc.). Lastly, the editing prompts that modify the colors (e.g., blue car to white car) can be quite effective in practice.

### v2

@@ 3.3 Ditail for Model-Centric Style Transfer
6:${\boldsymbol{c}}\leftarrow\alpha\,{\boldsymbol{e}}_{pos}-\beta\,{\boldsymbol{e}}_{neg}$triangleright Control the strength of content
@@
In Line 2-3 of Alg. 2, we first generate a guidance condition vector ${\boldsymbol{c}}$, an interpolation between the positive and negative prompt embedding using two hyper-parameters $\alpha,\beta$. $\alpha$ and $\beta$ are extra knobs to control the strength of the guidance for content injection. As shown in Fig 3(a), a larger $\alpha$ value increases the impact of the source image content, mainly generated by interpreting the positive prompt. Conversely, a larger $\beta$ value does not always yield a significant outcome, as it depends on the semantics of the negative prompt. If the negative prompt is relevant to the content/structure, then $\beta$ will change the resulting content guidance accordingly.
@@
Figure 3: (a) Visualization of noisy latent after DM inversion on real and generated images with different scaling factor $\alpha$. A larger $\alpha$ leads to a stronger structure prior embedded in the noisy latent, resulting in a more structure-preserved starting point in the reverse sampling process; (b) Demonstration of style transfer across different DMs. The diagonal images are the non-modified images generated using each diffusion model with the same textual prompt. Every column corresponds to the images generated by one diffusion model, using contents from other models. Better view with color. Zoom in for best view.
@@
We have two extra inputs for applying Ditail on the model-centric style transfer task. Firstly, the user may provide positive and negative prompts to describe the image to help the inversion process. In Appendix, we have also tested empty prompts, prompts generated from image captioning, and irrelevant prompts, which verify that both empty and captioning prompts can generate promising results. Secondly, the user needs to specify a DM for diffusion inversion. Ideally, a DM matching the content image domain is desirable. Empirically, the dependence on the inversion model is not strict, as we found that the generation results using different DMs to invert real content images are similar. Besides, Ditail adopts the invert-and-denoise workflow for the following two reasons: 1) the generating DM or the latent representations are not available in most cases, and 2) we want to adjust the condition scaling parameters $\alpha,\beta$ to strengthen or weaken the content transfer.
@@
We note this is a novel style transfer setting as we use DMs to set the domains/styles of interest. This allows us to benefit from the vast collection of fine-tuned DMs with drastically different specialization. If a certain style not included in the available DMs, we can merge existing DM checkpoints to create a novel target style (Sec. 3.5) or fine-tune a new DM model. Compared to the classic style transfer task, this is equivalent to defining the target style using a set of images. On the one hand, this gives us more flexibility in defining desirable styles. On the other hand, the set of images tends to provide more robust and comprehensive style information compared to a single image.
3.4 Diverse Applications of Ditail
Based on the core workflow of Ditail (Alg. 2), we can easily extend it to enable diverse applications, such as novel image generation and content manipulation.
Novel Image Generation. Given a pair of DMs $\theta^{(1)}$ and $\theta^{(2)}$, we input the positive and negative prompts to generate two images with each DM independently (denoted as $\mathcal{I}^{(1)}$ and $\mathcal{I}^{(2)}$), where each of them inherits certain contents and styles from their generating DMs. By inverting the images to the noisy latents $\{{\boldsymbol{z}}_{t}^{(1)}\}_{t=1:T}$ and $\{{\boldsymbol{z}}_{t}^{(2)}\}_{t=1:T}$ with some condition scalers $(\alpha,\beta)$, we can then apply Ditail to generate two more images, $\mathcal{I}^{(1)\rightarrow(2)}$ and $\mathcal{I}^{(2)\rightarrow(1)}$, each consisting of the structure information of one DM and the style information of the other. Theoretically speaking, given a set of $m$ DMs, we can use Ditail to generate $\mathcal{O}(m^{2})$ images over various configurations for condition scaling and target style DMs/LoRAs. This dramatically expands the generation capacity as opposed to the $m$ images by running every DM independently, producing images that are not attainable through a single DM. It also allows the user to have better control over the generation process as well as the integration of more domain-specific DMs. The user may first use some DMs to generate specific content structure and then transfer the spatial information to other DMs with desirable styles. Fig. 3(b) shows the cross-transfer over three DMs.
Content Manipulation. As shown in Fig. 1, Ditail offers two ways of manipulating the content in the source images: prompt-based and collage-based. Firstly, image content can be directly modified through positive and negative prompts (more examples in Appendix). Secondly, one can create a collage that demonstrates the desirable content structure, such as adding a pair of sunglasses to a human face (Fig. 2(b)), then feed it to the Ditail algorithm (Alg. 2) with descriptive prompts and DM selections. Ditail is able to align the appearance of the collage to have higher fidelity. We note that the target domain/style can be identical to the source image, meaning that we can use the same photo-realistic DM for both inversion and sampling stages to get an image of “A man wearing sunglasses” in the photo-realistic domain.
Figure 4: Qualitative comparison of style transfer results on generated images drawn from GEMRec-18K [8]. The results of ControlNet [42] and PnP [39] are obtained with the target DM checkpoint. Better view with color. Zoom in for best view.
3.5 Tailoring Novel Target Domains for Ditail
Ditail enables flexible target domain tailoring on basis of two key techniques: LoRAs and checkpoint merging. LoRA [14] has emerged as a powerful parameter-efficient tuning technique. There is a vast community of fine-tuning LoRAs dedicating to various characters and styles for image generation, which essentially consists of low-rank approximated parameters for the text encoders and attention layers in U-Net. At inference time, such parameters are then scaled and added to a DM base model. For a single layer with weight matrix $W\in\mathcal{R}^{m\times n}$, we fuse the LoRA parameters $B\in\mathcal{R}^{m\times r}$ and $A\in\mathcal{R}^{r\times n}$ via $\hat{W}=W+\lambda BA$, where $\hat{W}$ is the fused parameters and $\lambda$ is a weight scaler. Alternatively, direct checkpoint merging is also a simple yet effective approach for mixing multiple DMs, which essentially computes a weighted sum over all parameters. Formally, to obtain an interpolated style DM $\theta^{{}^{\prime}}$ between style DMs $\theta^{(1)}$ and $\theta^{(2)}$, we let $\theta^{{}^{\prime}}=\gamma\,\theta^{(1)}+(1-\gamma)\,\theta^{(2)}$, where $\gamma\in[0,1]$ is a hyper-parameter that controls the interpolation of the two styles. We present some qualitative results of checkpoint interpolation in Appendix.
@@ 4.1 Results on Generated Images
Dataset and Model Source. Utilizing Ditial for style transfer requires DM checkpoints with various domain specializations. The GEMRec-18K [8] dataset contains 18 thousand images generated by 200 text-to-image DMs fine-tuned on Stable Diffusion [30], where the DM checkpoints were collected from Civitai. Each DM generates 90 images, and the prompts are divided into 12 categories. We pick five DM checkpoints with distinctive style specialization (i.e., abstract, anime, fantasy, illustration, and photo-realistic) and perform any-to-any style transfer using the Ditail framework. We use the images $\{\mathcal{I}^{(1)}_{i}\}_{i=1:90}$ generated by one source DM $\theta^{(1)}$ as the content image and stylize it with another target DM $\theta^{(2)}$, resulting in $\{\mathcal{I}^{(1)\rightarrow(2)}_{i}\}_{i=1:90}$.
Figure 5: Qualitative results and ablation studies on real images from COCO Captions [3]. Better view with color. Zoom in for best view.
@@
Qualitative Results. We compare the qualitative results with other style transfer [21, 43], image manipulation [39, 1, 27], and conditional generation [42] approaches in Fig. 4. The content and style images are generated by $\theta^{(1)}$ and $\theta^{(2)}$ on the same prompt, respectively. Notably, performing style transfer with Ditail does not require a style image due to its model-centric nature, whereas it is a necessity for conventional style transfer approaches. One key observation is that training with one or a few style examples does not guarantee capturing the target style. Such failure will cause a near-unmodified output image, such as the car example for InST [43]. AdaAttn [21] works in most cases, yet the scattered colors tend to blur the subject in most examples. It is undeniable that the term “style” has been loosely defined, and most conventional style transfer works tend to specialize in the styles elicited from artistic images. However, we want to expand our scope to a more generalizable setting, which involves domain adaptation capabilities such as “realistic $\rightarrow$ anime” and “watercolor $\rightarrow$ illustration”, or even the nuance style that a single style image or plain text guidance can hardly capture. In such cases, numerous DM checkpoints with stable styles can serve as a plug-and-play choice and exhibit reasonable performance. As for other approaches, Instruct Pix2Pix [1] and Pix2Pix-Zero [27] hardly work on this task, whereas ControlNet [42] and PnP [39] are not robust enough even with the target style DM. More specifically, PnP fails on row 1&2 and ControlNet fails on row 3&4. Moreover, Ditail is better on details (e.g., color and lighting) when depicting the same target style, such as row 5&6.
@@
Ablation Study on Content Injection and Condition Scaling. In this experiment, we study the effects of content injection and various choices of $\alpha,\beta$ values for condition scaling. To quantify the change in image fidelity, structure, and style, we use three metrics, respectively. We use the cosine similarity in CLIP-Score [10, 28] to measure the compliance between prompt and image; we use the DINO self-similarity loss [2, 38] to measure the preservation of structure, which essentially compares the self-attention maps between source and target images; we use Fréchet Inception Distance (FID) [11] to the images generated by the target model to measure the closeness of style. By the results in Tab. 1, we observe the following: First, all CLIP scores do not deviate too much from the source image baseline, and the transformed images from most groups move towards the distribution of target images. Second, both enlarging $\alpha$ and introducing $\beta$ leads to a better-preserved structure but pushes the transformed images further away from the target style. Third, the effect of introducing $\beta$ is not monotonic when enlarging $\alpha$, as the negative prompt is generally not related to content or style. Fourth, the best-stylized results are obtained when $\alpha=2$. Lastly, similar to the conclusions from qualitative results, applying content injection alone already provides a strong baseline for structure preservation.
[…]
$\alpha=1,\beta=0$ 0.2586 0.0835 175.94
$\alpha=2,\beta=0$ 0.2591 0.0738 176.02
$\alpha=4,\beta=0$ 0.2606 0.0490 190.03
$\alpha=8,\beta=0$ 0.2560 0.0341 216.02
$\alpha=1,\beta=1$ 0.2589 0.0549 189.20
$\alpha=2,\beta=1$ 0.2579 0.0624 183.26
$\alpha=4,\beta=1$ 0.2590 0.0547 185.10
$\alpha=8,\beta=1$ 0.2562 0.0352 213.70
$\alpha=1,\beta=0$ 0.2546 0.0401 200.08
$\alpha=2,\beta=0$ 0.2585 0.0338 196.38
$\alpha=4,\beta=0$ 0.2575 0.0261 211.11
$\alpha=8,\beta=0$ 0.2538 0.0241 228.48
$\alpha=1,\beta=1$ 0.2563 0.0280 206.82
$\alpha=2,\beta=1$ 0.2566 0.0304 199.92
$\alpha=4,\beta=1$ 0.2574 0.0272 204.96
$\alpha=8,\beta=1$ 0.2537 0.0245 227.55
Table 1: Quantitative results for model-centric style transfer on generated images from the GEMRec-18K [8] dataset. We manually pick 5 models with distinctive styles and conduct style transfer on 90 textual prompts. The metrics are averaged over all 20 possible style transfer combinations. The mean CLIP cosine similarity [10, 28] on source images is 0.2587, and the mean FID [11] distance between source and target images is 227.32. The best-performed metric for each group is shown in bold.
@@ 4.2 Results on Real Images
Dataset and Model Source. We randomly sample 300 images from the COCO Captions [3] dataset. We use the same set of style DMs as in the previous experiments.
Qualitative Results. As shown in Fig. 5, we perform style transfer on the selected DMs with ablation studies on condition scaling and content injection. We use a unified negative prompt “worst quality, blurry, NSFW” by default. It can be seen that solely using scaling factors or content injection can preserve the content structure to some extent, and applying content injection, even with reconstructed features and self-attention maps, gives a remarkable structure-preserving performance. Besides, enlarging $\alpha$ leads to better structure preservation effect, which complies with the intuition provided by Fig. 3(a). In general, applying both techniques together enables us to keep the content unchanged while shifting to the target style seamlessly.
Figure 6: (a) Comparison of DINO (structure) and FID (style) with PnP [39] and ControlNet [42]; (b) Comparison of conditional scaling variants on structure preservation; zoom-in for best view.
Quantitative Comparison with Other Methods. We make quantitative comparisons of DINO self-similarity scores (structure preservation) and FID scores (style) with the best competitors, PnP [39] and ControlNet [42]. We consider three variants: prepending style keywords to the textual prompt (on a generic DM), using with the DM checkpoint of target style, and using both of them. As for Ditail, we perform inversion with a photo-realistic DM and then perform noise sampling with the target style DM. We use the condition scaling configurations with the space of $\alpha\in\{1,2,4,8\}$ and $\beta\in\{0,1\}$. By taking a closer look at Fig. 6(a), we draw the following conclusions: 1) PnP and ControlNet with pre-pending text gives poor stylization performance, suggesting that the nuance styles can be hardly expressed via plain text signals; 2) PnP and ControlNet can exhibit reasonable performance using the target style DM, but may still suffer from robustness issues as shown in Fig. 4, which also demonstrates the superiority of domain-specific DMs compared to generic ones; 3) Ditail enables a flexible trade-off between on structure preservation (DINO) and stylization (FID) via the condition scaling factors $\alpha$ and $\beta$, resulting in better output images in both dimensions.
Why Condition Scaling Works. To get a better intuition for the condition scaling technique, we compare the proposed version (i.e., direct weighted sum in the Euclidean space, ‘raw’) with two other variants with normalization. Specifically, ‘norm’ normalizes embedding vector ${\boldsymbol{c}}$ after the weighted sum, whereas ‘scaled norm’ rescale the normalized vector by the mean norm. We evaluate the DINO scores with $\alpha\in\{1,2,4,8\}$ and $\beta=1$. From Fig. 6(b) we can observe that: 1) a wrong condition direction can lead to structure collapse (e.g., $\alpha=1$ in cases with normalization); 2) normalized condition vectors tend to exhibit similar and weaker structure preservation effect; 3) rescaling the normalized condition vectors restores the content to some extent. Hence, the magnitude of condition vectors seem to be more important in maintaining the content consistency. Our hypothesis is that a larger magnitude of the condition vector can interfere the inversion process via the reversely predicted noise, resulting in less random noise being added to the latents and thereby preserving the content. Besides, apart from being a flexible structure preservation technique, and we also empirically found that condition scaling can be helpful in bridging the distribution gap of hidden states from different DM checkpoints.
5 Limitation and Social Impact
Despite the diverse side applications empowered by Ditail, it still suffers from a few disadvantages. First, Ditail involves an extra DDIM [36] inversion process, thus resulting in a higher latency compared to conventional text-to-image generations. This can potentially be mitigated via LCMs [22, 23], which substantially reduces the number of inference steps. Second, Ditail may not generalize to mix DMs with distinctive architectures, as the content injection expects hidden states of identical dimensions. Third, Ditail has a strong dependency on checkpoint quality (e.g., fidelity and robustness on the target domain), so the generation results may not be ideal with target DMs of poor quality. Last, Ditail has limited cooperation with human annotations due to training-free nature of the method.
Meanwhile, we also want to emphasize some potential impact elicited by Ditail. First, due to the probabilistic nature of generative models, the images they generate may inadvertently contain potentially unsafe content. Thus, it is essential to adopt content safety checkers in production environment, ensuring that any inappropriate content is filtered before dissemination. Second, as community uploaders fine-tune DMs with private datasets, there are potential copyright concerns regarding the training data and the checkpoint itself. Moreover, the copyright issue also exists for the merged checkpoints with tailored target styles, where the copyright attribution can be tough to determine even with DM fingerprinting techniques. We stay alert to these concerns and maintain ongoing vigilance in monitoring them closely. Last, Ditail’s strong capability of diversified image generation can boost the artistic creation of non-academic users on existing DM/LoRA platforms and may potentially lead to business benefits.